OpenStudy (anonymous):

tan^2a+3=3seca find a?

OpenStudy (anonymous):

you have tan^2X= sec^2X -1 subtitute it into the eqn, we have (sec^2a -1) + 3 =3seca <=> sec^2a - 1 +3 = 3seca <=> sec^2a +2 -3seca = 0 <=> seca= 2 or seca = 1 seca=1/cosa so 1/cosa=2 or 1/cosa=1 therefore cosa = 0.5 or 1 so a = 60 or 360, and depends on the range of answers they allowed, just adjust your ans using the astc circle